# Optimizing Multi-modality Trackers via Sensitivity-regularized Tuning

### v2

@@ IV-C Ablation Study and Analysis
Event Frame [38] 66.8 95.5 61.2 77.6 68.8 82.0
Event Count [33] 66.3 95.1 59.8 76.9 67.9 81.2
Time Surface [39] 66.1 94.9 60.3 77.1 67.4 80.6
TSLTD [6] 66.5 95.2 61.0 77.4 67.9 81.3
Event Volume [69] 67.3 96.1 61.4 77.9 69.1 82.4
@@
Effectiveness of Smaller Learning Rates. In this paper, we exploit the sensitivity-regularized gradients to bolster the transfer process. A straightforward strategy might involve using a smaller learning rate. To investigate this, we evaluate the effect of reduced learning rates. As shown in Table XI, a smaller learning rate (i.e., $\alpha=10^{-5}$) yields negligible improvement in PR, but at the cost of a decline in SR. Further reduction (i.e., $\alpha=10^{-6}$) leads to overall performance degradation. These findings suggest that merely reducing the learning rate is insufficient to address over-fitting in cross-modality adaptation, while simultaneously limiting transfer potential, as it uniformly suppresses parameter updates without prioritizing those that are sensitive or high-risk.
TABLE XI: Ablation analysis of smaller learning rates $\alpha$ on the LasHeR dataset.
@@
Observations on Weight Variations. To intuitively demonstrate the effectiveness of the proposed sensitivity-aware regularization tuning, we visualize parameter dynamics during training from both spatial and temporal perspectives, as depicted in Fig. 9 and Fig. 10. Overall, our regularization encourages smaller and more evenly distributed weight deviations, effectively mitigating over-fitting and facilitating stable adaptation. As shown in Fig. 9, for parameters with high prior sensitivity, such as blocks.3.attn.proj.weight (highlighted in the orange box), our method effectively suppresses excessive updates. As further demonstrated in Fig. 10, our method facilitates faster convergence and achieves a substantial reduction in sensitivity-weighted weight deviations (e.g., up to $80\%$), indicating improved retention of the pre-trained knowledge while ensuring desired task adaptation. Moreover, the vanilla fine-tuning causes parameter updates to concentrate excessively in localized regions, hindering the global adaptability. After applying the sensitivity-aware penalties, the parameter updates become more balanced and distributed, effectively decentralizing the adaptation burden from a few parameters to a broader set. This redistribution helps mitigate adaptation risk and promotes more stable fine-tuning.
@@
Figure 10: Sensitivity-weighted weight distances between our regularized tuning method and the full fine-tuning (with different learning rates) on the DepthTrack.
TABLE XII: Computational cost of Prior Sensitivity Estimation on OSTrack-B256.
Metric Computational Cost
Operation-wise Model-wise Total Process
FLOPs 104.8 G 15.5 T $\sim$ 39.7 E
Latency $294.5$ ms $43.0$ s 47.8 h ($\sim$2 day)
Memory $164.2$ GB - -
TABLE XIII: Computational complexity and speed analysis on the LasHeR dataset.
Method Base Model Param (M) FLOPs (G) FPS SR PR
OSTrack [58] OSTrack 92.1 58.1 98.4 47.2 58.5
ProTrack [57] OSTrack 92.7 58.4 92.3 42.0 53.8
ViPT [70] OSTrack 92.9 59.9 88.6 52.5 65.1
SDSTrack [17] OSTrack 102.1 108.7 44.6 53.1 66.5
UnTrack [51] OSTrack 98.7 62.6 75.6 51.3 63.7
ViPT+Ours OSTrack 92.9 59.9 88.6 54.9 68.4
UnTrack+Ours OSTrack 98.7 62.6 75.6 54.0 67.4
Ours OSTrack 202.0 149.0 49.0 56.3 70.1
DropTrack [49] DropTrack 92.1 130.6 57.6 47.7 59.2
ViPT [70] DropTrack 92.9 131.9 48.8 52.5 65.6
SDSTrack [17] DropTrack 102.1 244.2 26.8 55.1 69.0
UnTrack [51] DropTrack 98.7 138.2 42.3 52.4 65.9
ViPT+Ours DropTrack 92.9 131.9 48.8 57.1 70.5
UnTrack+Ours DropTrack 98.7 138.2 42.3 56.3 69.4
Ours DropTrack 202.0 335.3 29.2 58.8 73.6
SUTrack [9] SUTrack 65.29 106.9 37.6 60.8 75.6
Ours SUTrack 141.1 243.7 24.8 62.9 77.8
IV-D Computational Efficiency
@@
In addition to tracking accuracy, we evaluate the computational efficiency of our sensitivity-aware regularization tuning by analyzing both training and inference overhead.
Computational Overhead of Prior Sensitivity. Prior sensitivity is estimated via a group-wise FIM approximation with eigen-decomposition. With the operation-diagonal (group-wise) approximation and Rayleigh-quotient probing, we reduce the computational cost from intractable to a practical level. We further profile this procedure on a cluster of 8 NVIDIA RTX 3090 Ti GPUs, reporting theoretical FLOPs, latency, and peak GPU memory in Table XIV to quantify its efficiency. While this cost is justified by the resulting performance gains, we explicitly acknowledge that—unlike plug-and-play PEFT methods—prior sensitivity estimation introduces additional, non-trivial computation. Importantly, it is a one-shot, offline pre-processing step performed only once before fine-tuning, and thus incurs negligible overhead during subsequent training iterations and inference. Moreover, it can accelerate convergence, effectively offsetting the preprocessing time and potentially reducing the overall time-to-result across the full training lifecycle (see Appendix for details).
@@
Efficiency of Sensitivity-Regularized Fine-tuning. We compare the training efficiency of our method against vanilla fine-tuning. Since prior sensitivity estimation is conducted as a one-time preprocessing step before fine-tuning, its runtime is excluded from the training time. Our method introduces parameter-wise transfer sensitivity quantification with marginal computational overhead, as this sensitivity is derived from the off-the-shelf gradients. While vanilla fine-tuning operates at 37.5 ms per iteration, our method runs at 39.0 ms per iteration, incurring only a modest $4.0\%$ increase in training time.
@@
Inference Cost and Speed. We compare the computational complexity and runtime efficiency of representative baselines, focusing on trackers instantiated with three widely used pre-trained backbones: OSTrack, DropTrack and SUTrack. For a fair comparison, we report both theoretical cost (e.g., parameter count and FLOPs) and practical throughput (FPS) under a unified evaluation protocol. All experiments are conducted on the same hardware platform, equipped with an Intel(R) Xeon(R) 6456C CPU, 256 GB RAM, and a single NVIDIA RTX 3090Ti GPU. As reported in Table XIII, our method supports real-time tracking with 24.8 FPS while delivering top-tier accuracy under the SUTrack setting. When integrated into ViPT (“ViPT+Ours”), it yields both superior speed and performance. Under the DropTrack base model, the resulting tracker still operates at 48.8 FPS, maintaining a respectable speed given its focus on accuracy.
IV-E Limitations and Discussion
@@
From a computational-efficiency perspective, a practical consideration is that prior sensitivity estimation (group-wise FIM approximation with Rayleigh-quotient probing) can still be a non-trivial extra offline procedure compared with plug-and-play PEFT: even after the proposed approximations make it practical, it can still be relatively heavyweight, which can become a hard barrier for large backbones, frequent architecture changes, or rapid ablation cycles. Additionally, the formulation defines the Fisher expectation over the pre-training dataset, so applying the method as-is may require access to (or a governed surrogate for) pre-training data—an assumption that does not always hold under licensing, privacy, or retention constraints—potentially limiting portability across model sources and deployment settings. Addressing these constraints, e.g., by exploring data-free or proxy-based sensitivity estimation—remains an important direction for future work.
V Conclusion
@@
This study revisits the critical misfitting issues encountered when adapting pre-trained RGB trackers to multi-modal tracking tasks. By introducing two complementary sensitivities that capture the dynamic shift in parameter importance as models transition from pre-trained to multi-modal contexts, our regularized tuning method strategically calibrates gradient updates. Extensive experiments demonstrate the effectiveness of our method, surpassing current state-of-the-art techniques across various multi-modal tracking scenarios, with notable performance improvements resulting from the incorporation of regularization. A key insight from our work is the importance of sensitivity-aware fine-tuning, which offers superior generalization and adaptability compared to traditional fine-tuning methods. Ultimately, our findings highlight the need for more nuanced cross-domain generalization strategies, as overly rigid or flexible fine-tuning can hinder pre-trained model transferability. We believe these insights will pave the way for further advancements in multi-modal transfer learning, particularly in the context of scene perception tasks.
Appendix A Proof of Derivations
A-A Proof of Proposition 1
@@
  1. Bounded FIM Error. Each group approximation $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}$ captures the principal tangent of $\mathcal{F}^{(\theta_{0}^{j})}$ with bounded error. Due to symmetry and orthogonal eigenvectors, the Frobenius norm of the error is derived from the difference in their eigenvalues:
@@
  2. Bounded Generalization Gap Error. The approximate FIM $\tilde{\mathcal{F}}^{(\theta_{0})}$ remains close to $\mathcal{F}^{(\theta_{0})}$ as a metric on parameter space. For any parameter difference $\Delta\theta\in\mathbb{R}^{|\theta|}$, let $\varepsilon_{gen}(\mathcal{F}^{(\theta_{0})})=\frac{1}{2}\Delta\theta^{T}\mathcal{F}^{(\theta_{0})}\Delta\theta$ denote the generalization gap induced by the true FIM. Likewise $\varepsilon_{gen}(\tilde{\mathcal{F}}^{(\theta_{0})})=\frac{1}{2}\Delta\theta^{T}\tilde{\mathcal{F}}^{(\theta_{0})}\Delta\theta$ is the approximated generalization gap. Then the discrepancy between these distances is bounded in terms of the residual generalization gap error. In particular, one has following formula by definition of the spectral norm:
@@
A-B Proof of Transfer Sensitivity
We state the local assumptions under which the transfer sensitivity bound in Eq. (9)-(12) holds and provide a more explicit derivation.

### v3

@@ 4.3 Ablation Study and Analysis
Event Frame Rebecq et al. (2017) 66.8 95.5 61.2 77.6 68.8 82.0
Event Count Maqueda et al. (2018) 66.3 95.1 59.8 76.9 67.9 81.2
Time Surface Sironi et al. (2018) 66.1 94.9 60.3 77.1 67.4 80.6
TSLTD Chen et al. (2020) 66.5 95.2 61.0 77.4 67.9 81.3
Event Volume Zhu et al. (2019) 67.3 96.1 61.4 77.9 69.1 82.4
@@
Table 11: Ablation analysis of smaller learning rates $\alpha$ on the LasHeR dataset.
@@
Effectiveness of Smaller Learning Rates. In this paper, we exploit the significance-regularized gradients to bolster the transfer process. A straightforward strategy might involve using a smaller learning rate. To investigate this, we evaluate the effect of reduced learning rates. As shown in Table 11, a smaller learning rate (i.e., $\alpha=10^{-5}$) yields negligible improvement in PR, but at the cost of a decline in SR. Further reduction (i.e., $\alpha=10^{-6}$) leads to overall performance degradation. These findings suggest that merely reducing the learning rate is insufficient to address over-fitting in cross-modality adaptation, while simultaneously limiting transfer potential, as it uniformly suppresses parameter updates without prioritizing those that are sensitive or high-risk.
@@
Figure 10: Significance-weighted weight distances between our regularized tuning method and the full fine-tuning (with different learning rates) on the DepthTrack.
Observations on Weight Variations. To intuitively demonstrate the effectiveness of the proposed significance-aware regularization tuning, we visualize parameter dynamics during training from both spatial and temporal perspectives, as depicted in Fig. 9 and Fig. 10. Overall, our regularization encourages smaller and more evenly distributed weight deviations, effectively mitigating over-fitting and facilitating stable adaptation. As shown in Fig. 9, for parameters with high prior significance, such as blocks.3.attn.proj.weight (highlighted in the orange box), our method effectively suppresses excessive updates. As further demonstrated in Fig. 10, our method facilitates faster convergence and achieves a substantial reduction in significance-weighted weight deviations (e.g., up to $80\%$), indicating improved retention of the pre-trained knowledge while ensuring desired task adaptation. Moreover, the vanilla fine-tuning causes parameter updates to concentrate excessively in localized regions, hindering the global adaptability. After applying the significance-aware penalties, the parameter updates become more balanced and distributed, effectively decentralizing the adaptation burden from a few parameters to a broader set. This redistribution helps mitigate adaptation risk and promotes more stable fine-tuning.
Table 12: Computational complexity and speed analysis on the LasHeR dataset. The base trackers are denoted by abbreviations (e.g., OS/Drop/SU). Model size is reported in millions of parameters (M).
Method Model Params GFLOPs FPS SR PR
OSTrack Ye et al. (2022) OS 92.1 58.1 98.4 47.2 58.5
ProTrack Yang et al. (2022) OS 92.7 58.4 92.3 42.0 53.8
ViPT Zhu et al. (2023a) OS 92.9 59.9 88.6 52.5 65.1
SDSTrack Hou et al. (2024) OS 102.1 108.7 44.6 53.1 66.5
UnTrack Wu et al. (2024b) OS 98.7 62.6 75.6 51.3 63.7
ViPT+Ours OS 92.9 59.9 88.6 54.9 68.4
UnTrack+Ours OS 98.7 62.6 75.6 54.0 67.4
Ours OS 202.0 149.0 49.0 56.3 70.1
DropTrack Wu et al. (2023) Drop 92.1 130.6 57.6 47.7 59.2
ViPT Zhu et al. (2023a) Drop 92.9 131.9 48.8 52.5 65.6
SDSTrack Hou et al. (2024) Drop 102.1 244.2 26.8 55.1 69.0
UnTrack Wu et al. (2024b) Drop 98.7 138.2 42.3 52.4 65.9
ViPT+Ours Drop 92.9 131.9 48.8 57.1 70.5
UnTrack+Ours Drop 98.7 138.2 42.3 56.3 69.4
Ours Drop 202.0 335.3 29.2 58.8 73.6
SUTrack Chen et al. (2025b) SU 65.29 106.9 37.6 60.8 75.6
Ours SU 141.1 243.7 24.8 62.9 77.8
4.4 Computational Efficiency
@@
In addition to tracking accuracy, we evaluate the computational efficiency of our significance-aware regularization tuning by analyzing both training and inference overhead.
Computational Overhead of Prior Significance. Our method requires an offline estimation of prior significance (e.g., FIM-based eigen-decomposition and Rayleigh quotient probing), which introduces a non-trivial preprocessing cost compared to purely plug-and-play PEFT baselines. This overhead can become a practical barrier when scaling to larger backbones or when frequent architectural changes demand repeated recomputation. Nonetheless, this cost is paid once and does not materially affect the per-iteration training and inference runtime after the significance is computed (details in the Appendix C).
@@
Efficiency of Significance-Regularized Fine-tuning. We compare the training efficiency of our SRFT against vanilla FFT. Since prior significance is computed once as a preprocessing step, we exclude it from per-iteration training time. SRFT introduces parameter-wise transfer significance quantification with marginal computational overhead, as this significance is derived from the off-the-shelf gradients: 39.0 ms/iter vs. 37.5 ms/iter for vanilla FFT ($4.0\%$ ).
@@
Inference Cost and Speed. We compare the computational complexity and runtime efficiency of representative baselines, focusing on trackers instantiated with three widely used pre-trained backbones: OSTrack, DropTrack and SUTrack. For a fair comparison, we report both theoretical cost (e.g., parameter count and FLOPs) and practical throughput (FPS) under a unified evaluation protocol. All experiments are conducted on the same hardware platform, equipped with an Intel(R) Xeon(R) 6456C CPU, 256 GB RAM, and a single NVIDIA RTX 3090Ti GPU. As reported in Table 12, our method supports real-time tracking with 24.8 FPS while delivering top-tier accuracy under the SUTrack setting. When integrated into ViPT (“ViPT+Ours”), it yields both superior speed and performance. Under the DropTrack base model, the resulting tracker still operates at 48.8 FPS, maintaining a respectable speed given its focus on accuracy.
4.5 Limitations and Discussion
@@
From a computational-efficiency perspective, a practical consideration is that prior significance estimation (group-wise FIM approximation with Rayleigh-quotient probing) can still be a non-trivial extra offline procedure compared with plug-and-play PEFT: even after the proposed approximations make it practical, it can still be relatively heavyweight, which can become a hard barrier for large backbones, frequent architecture changes, or rapid ablation cycles. Additionally, the formulation defines the Fisher expectation over the pre-training dataset, so applying the method as-is may require access to (or a governed surrogate for) pre-training data—an assumption that does not always hold under licensing, privacy, or retention constraints—potentially limiting portability across model sources and deployment settings. Addressing these constraints, e.g., by exploring data-free or proxy-based significance estimation—remains an important direction for future work.
5 Conclusion
@@
This study revisits the critical misfitting issues encountered when adapting pre-trained RGB trackers to multi-modal tracking tasks. By introducing two complementary parameter significance that capture the dynamic shift in parameter importance as models transition from pre-trained to multi-modal contexts, our regularized tuning method strategically calibrates gradient updates. Extensive experiments demonstrate the effectiveness of our method, surpassing current state-of-the-art techniques across various multi-modal tracking scenarios, with notable performance improvements resulting from the incorporation of regularization. A key insight from our work is the importance of significance-aware fine-tuning, which offers superior generalization and adaptability compared to traditional fine-tuning methods. Ultimately, our findings highlight the need for more nuanced cross-domain generalization strategies, as overly rigid or flexible fine-tuning can hinder pre-trained model transferability. We believe these insights will pave the way for further advancements in multi-modal transfer learning, particularly in the context of scene perception tasks.
Data Availability
This work does not propose a new dataset. All the datasets we used are publicly available and well cited in the paper. FE108: https://zhangjiqing.com/dataset/; VisEvent: https://sites.google.com/view/viseventtrack/; CoeSot: https://github.com/Event-AHU/COESOT; DepthTrack: https://github.com/xiaozai/det; VOT RGBD2022: https://www.votchallenge.net/vot2022/dataset.html; LasHeR: https://github.com/BUGPLEASEOUT/LasHeR; RGBT234: https://github.com/xuboyue1999/RGBT-Tracking/tree/main.
Conflict of Interest
The authors affirm that there are no commercial or associative relationships that could be perceived as a conflict of interest related to the submitted work.
Appendix
This appendix contains the following contents. In Section A, we provide implementation details for multi-modal tracking, including a more detailed description of the network architecture. In Section B, we present the complete derivations and proof details. In Section C, we report additional quantitative results, including comparisons with SPT, the computational overhead introduced by prior significance estimation, the training efficiency of SRFT, and different significance fusion scheduling strategies. Finally, we have supplemented some tracking visuals for a better qualitative comparison in Section D.
  • Section A: Implementation details.
  • Section B: Proof of derivations.
  • Section C: Additional ablation studies.
  • Section D: Visualization of tracking results.
Appendix A Implementation details
The detailed structure of the multimodal tracker is shown in Fig. 11. The input of our proposed tracking model consists of a pair of template frames and a pair of search frames, i.e., one RGB template frame $\mathrm{Z}_{\mathrm{R}}\in\mathbb{R}^{H_{z}\times W_{z}\times 3}$, one RGB search frame $\mathrm{X}_{\mathrm{R}}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}$, one auxiliary-modal template frame $\mathrm{Z}_{\mathrm{A}}\in\mathbb{R}^{H_{z}\times W_{z}\times 3}$, and one auxiliary-modality search frame $\mathrm{X}_{\mathrm{A}}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}$. Notably, to make event data compatible with the RGB domain, we aggregate the event set between the image and its next one into a three-channel color-polar event image. These data are first split and flattened into sequences of patches $\mathrm{z}_{\mathrm{R}},\mathrm{z}_{\mathrm{A}}\in\mathbb{R}^{N_{z}\times\left(3P^{2}\right)}$ and $\mathrm{x}_{\mathrm{R}},\mathrm{x}_{\mathrm{A}}\in\mathbb{R}^{N_{x}\times\left(3P^{2}\right)}$, where $P\times P$ is the resolution of each patch ($P=16$), and $N_{z}=\frac{H_{z}W_{z}}{P^{2}}$, $N_{x}=\frac{H_{x}W_{x}}{P^{2}}$. Next, two modal-aware patch embedding layers are used to project $\mathrm{z}_{\mathrm{R}},\mathrm{x}_{\mathrm{R}}$ and $\mathrm{z}_{\mathrm{A}},\mathrm{x}_{\mathrm{A}}$ into the D-dimensional latent space, $\mathrm{z}_{\mathrm{R}},\mathrm{z}_{\mathrm{A}}\in\mathbb{R}^{N_{z}\times D}$ and $\mathrm{x}_{\mathrm{R}},\mathrm{x}_{\mathrm{A}}\in\mathbb{R}^{N_{x}\times D}$. The patch embeddings $\mathrm{z}_{\mathrm{R}}$ and $\mathrm{x}_{\mathrm{R}}$ are concatenated as $\mathbf{H}_{\mathrm{R}}^{(0)}=\left[\mathrm{z}_{\mathrm{R}};\mathrm{x}_{\mathrm{R}}\right]\in\mathbb{R}^{\left(N_{z}+N_{x}\right)\times D}$, and $\mathrm{z}_{\mathrm{A}}$ and $\mathrm{x}_{\mathrm{A}}$ are concatenated as $\mathbf{H}_{\mathrm{A}}^{(0)}=\left[\mathrm{z}_{\mathrm{A}};\mathrm{x}_{\mathrm{A}}\right]\in\mathbb{R}^{\left(N_{z}+N_{x}\right)\times D}$. The computation of modality-aware ViT block can be formulated as:
$\begin{gathered}\mathbf{H}_{\mathrm{X}}^{\prime(l)}=\mathbf{H}_{\mathrm{X}}^{(l-1)}+\operatorname{MSA}\left(\mathrm{LN}\left(\mathbf{H}_{\mathrm{X}}^{(l-1)}\right)\right)\\ \mathbf{H}_{\mathrm{X}}^{(l)}=\mathbf{H}_{\mathrm{X}}^{\prime(l)}+\operatorname{MLP}\left(\mathrm{LN}\left(\mathbf{H}_{\mathrm{X}}^{\prime(l)}\right)\right)\end{gathered}$
[…]
Figure 11: More detailed network architecture of our multi-modality trackers.
Appendix B Proof of Derivations
B.1 Validity of Loss-Parameter Manifold Hypothesis
The loss-parameter manifold hypothesis is the mathematical foundation of modern deep learning optimization. It posits that, for a given model and dataset, the loss function induces a highly structured geometric object—a manifold (or manifold-like set)—embedded in parameter space, and that training dynamics can be understood as motion on or near this surface. For the manifold hypothesis, a large body of prior work (Bengio et al., 2013; Pope et al., 2021; Song and Ermon, 2019; Kiani et al., 2024) has widely and successfully employed manifold-based assumptions to analyze representation learning and generalization behavior in deep networks. While these works consistently demonstrate that the manifold assumption may not hold globally, local manifold structure provides a meaningful and empirically validated abstraction for reasoning about model behavior under realistic perturbations—precisely the regime considered in our analysis. In this view, parameters with similar loss values form smooth level sets, locally approximable by tangent planes, while curvature governs how sensitive the loss is to perturbations in different directions. In this work, the prior significance does not assume that a global manifold hypothesis holds in multi-modality tracking scenarios. Instead, our method is deliberately local: it characterizes the behavior of the pretrained objective in a neighborhood of the pretrained parameters $\theta_{0}$ with pre-trained dataset $D_{0}$, where our prior significance estimation operates. As discussed in Section 3, the prior significance approximation defines a practical and computable measure of the increase in pretrained loss induced by parameter deviations around $\theta_{0}$ under $D_{0}$, where the model gradient vanishes; this increase is characterized by the second-order Fisher information matrix (FIM). This formulation serves as a geometric metric, identifying the directions in parameter space most crucial for preserving pretrained knowledge, rather than making a claim about global parameter geometry. To ensure practical applicability, we leverage the empirically observed spectral concentration of the FIM in deep networks and adopt a low-rank approximation via Rayleigh-quotient probing, supported by an approximation error analysis.
B.2 Proof of Proposition 1
  1. 1.
@@ B.2 Proof of Proposition 1
    Bounded FIM Error. Each group approximation $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}$ captures the principal tangent of $\mathcal{F}^{(\theta_{0}^{j})}$ with bounded error. Due to symmetry and orthogonal eigenvectors, the Frobenius norm of the error is derived from the difference in their eigenvalues:
@@
  2. 2.
    Bounded Generalization Gap Error. The approximate FIM $\tilde{\mathcal{F}}^{(\theta_{0})}$ remains close to $\mathcal{F}^{(\theta_{0})}$ as a metric on parameter space. For any parameter difference $\Delta\theta\in\mathbb{R}^{|\theta|}$, let $\varepsilon_{gen}(\mathcal{F}^{(\theta_{0})})=\frac{1}{2}\Delta\theta^{T}\mathcal{F}^{(\theta_{0})}\Delta\theta$ denote the generalization gap induced by the true FIM. Likewise $\varepsilon_{gen}(\tilde{\mathcal{F}}^{(\theta_{0})})=\frac{1}{2}\Delta\theta^{T}\tilde{\mathcal{F}}^{(\theta_{0})}\Delta\theta$ is the approximated generalization gap. Then the discrepancy between these distances is bounded in terms of the residual generalization gap error. In particular, one has following formula by definition of the spectral norm:
@@
B.3 Derivation of Transfer Significance
We state the local assumptions under which the transfer parameter significance bound in Eq. (9)-(12) holds and provide a more explicit derivation.